OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the measure of B. 22° 38° 142° 158°

OpenStudy (anonymous):

@perl

OpenStudy (perl):

this is better http://www.mathwarehouse.com/triangle-calculator/online.php

OpenStudy (anonymous):

but itsn't that cheating?

OpenStudy (perl):

its a way to check ?

OpenStudy (anonymous):

no u can use calculators at school

OpenStudy (anonymous):

so its 38?

OpenStudy (perl):

side A is opposite angle A

OpenStudy (perl):

yes

OpenStudy (anonymous):

thnx

OpenStudy (perl):

you can use law of cosines by the way to solve this problem

OpenStudy (anonymous):

ok thnx

OpenStudy (perl):

a = 3, b = 5, and c = 7. we want angle B , which is between sides a and c 3^2 + 7^2 - 2*3*7*cos(theta) = 5^2 solve for theta cos(theta) = (5^2 - 3^2 - 7^2)/(-2*3*7) theta = arcos(5^2 - 3^2 - 7^2)/(-2*3*7) theta = 38.2132